A fire hose ejects a stream of water at an angle of 50.0 above the horizontal. The water behaves like a projectile, what is the highest place on the building the water can reach if the truck is located 30 m away from the building



1
Expert's answer
2021-05-25T16:50:27-0400

Gives

"\theta=50°"

Range=30 meter

We know that

Range

"R=\frac{u^2sin2\theta}{g}"

Put value

"30=\frac{u^2sin100°}{9.8}"

"u^2=\frac{30\times9.8}{sin100°}=298.53m/sec"

"u=17.28m\/sec"

Maximum highest place building of the water Hmax

"H=\frac{u^2sin^2\theta}{2g}"

Put value


"H=\frac{{17.28}^2sin^2100}{2\times9.8}=15.0m"